geerky42 (geerky42):

If the odds of a certain baseball team winning a game is \(\dfrac{4}{9}\) and there are 156 games in the regular season, then how many games can they except to win?

OpenStudy (anonymous):

odds are 4:9 meaning the probability of winning is \(\frac{4}{4+9}=\frac{4}{13}\)

geerky42 (geerky42):

why 4+9?

OpenStudy (anonymous):

they therefore expect to win \(\frac{4}{13}\times 156\)

OpenStudy (anonymous):

@satellite73 can u help me after?

OpenStudy (anonymous):

odd of winning is probability of winning divided by probability of losing another way to think about it is , if the odds are 4:9 then out of 13 games you win 4, lose 9